OpenStudy (anonymous):

A quartic function in which the coefficients are relatively prime is known to have zeros of 1+3i and -2 - √5 What is the coefficient of the cubic term of this function?

OpenStudy (anonymous):

Is there a quicker way than expanding this?

OpenStudy (anonymous):

You don't have to solve! Just tell me if there's a quicker way!

OpenStudy (btaylor):

The other zeroes are 1-3i and -2 + √5. I don't know about the -2 +/- √5, but I have a shortcut for the imaginary root. The two imaginary roots will multiply to form a trinomial. In the trinomial (ax^2 + bx + c), -b/a = (1-3i)+(1+3i) c/a=(1-3i)(1+3i). a is almost always 1. IDK, it may work for the squares. It should. Try it!

OpenStudy (anonymous):

Thank you! I never thought of doing it that way! I find the two quadratics and multiply to get the cubic coefficient! :)

OpenStudy (btaylor):

yup.